OpenStudy (anonymous):

What is the solution of y − 4x = 0 and 3x + 6y = 9?

OpenStudy (texaschic101):

do you prefer substitution or elimination

OpenStudy (anonymous):

substitution a) x=0, y=3/2 b) x=1/3, y=4/3 c) x=1 y=-1 d) x=1/4, y=2/3

OpenStudy (texaschic101):

y - 4x = 0 y = 4x now sub 4x in for y in the other equation 3x + 6y = 9 3x + 6(4x) = 9 3x + 24x = 9 27x = 9 x = 9/27 x = 1/3 now sub 1/3 in for x in either equation to find y y = 4x y = 4(1/3) y = 4/3 any questions ?